# Grounding Multimodal Large Language Models in Actions

### v1

@@ 1 Introduction
A key and open challenge in grounding MLLMs, which limits their capabilities in embodied tasks, is the gap between the native output space, natural language, and the action space of embodied agents. This problem is particularly acute in continuous action spaces, where low-level controllers may require a high degree of precision. Across the literature, a number of architectures and ways of handling action spaces have been proposed, but there has not been a systematic study of these designs. Our contributions generalize prior attempts to adapt MLLMs to generate actions through an empirical study on which principles and strategies are necessary to effectively close the gap between the action spaces of MLLMs and embodied agents. We study various grounding re-parameterization strategies, which we refer to as Action Space Adapter (ASA), across a range of embodiments, action spaces, and environments. In particular, we explore the following types of ASAs: (1) ASAs that directly generate actions from a new prediction policy using the MLLM hidden representations as input; (2) ASAs that reuse the native token space of the MLLM to encode actions; (3) and ASAs that introduce a new token space to encode the actions of the agent while adapting the MLLMs to predict these new tokens.
@@ 3.2.1 Discrete Action Spaces
Non-Semantic Language (Lang): Actions are mapped to language tokens, but instead of semantically meaningful descriptions of the actions as with SemLang, the actions are mapped to sequences of numbers. For example, “pick apple" is represented with the string “5 3". The policy must then output the tokens corresponding to this text to call this pick action.
@@ 3.2.2 Continuous Action Space Adaptors
where $v_{k}\in V$. The codebook $V$ of size $K$ is learned over an offline dataset $\mathcal{D}$ of actions using a VQ-VAE [54]. We overwrite $K$ infrequently used tokens from the LLM vocabulary to represent $V$. We defer the full details of this tokenization process to Section C.2.
@@
where $v_{k}^{i}\in V_{i}$ is the $k^{\textrm{th}}$ code from the $i^{\textrm{th}}$ codebook. Such tokenization can be learned using Residual VQ-VAE [RVQ-VAE, 50] on an offline dataset of actions. The actual number of token sequences we can represent is $K^{M}$. Hence, RVQ presents the opportunity to exponentially increase the action space quantization without having to drastically increase the size of the learned individual codebooks.
@@ 3.3 Training
We use LLaVA-1.5-7B [6] as the base MLLM. In the complex environments we consider, the zero-shot performance of MLLM is poor across all ASAs, even with detailed prompts. We therefore finetune the MLLM with interactive (i.e., action-labeled) data to make it more suited for interacting with a dynamic environment.
@@
Reinforcement Learning (RL) from Environment Feedback We can also optionally finetune the MLLM to optimize an environment reward using RL. However, predicting actions in the MLLM token space dramatically increases the number of possible action predictions, with many possible predictions corresponding to no valid action. For example, there are 32,000 tokens in the LLaMA text tokenizer, giving $32,000^{m}$ possible predictions by the model with $m$ tokens per action. This makes exploration difficult in RL as only a small fraction of the possible actions are valid. AMLM therefore uses a token filter to restrict the autoregressive sampling to only be from token sequences corresponding to valid actions. The token filter is a function $M(l_{t}^{1},\dots,l_{t}^{j-1})$ that produces a binary mask over all tokens to represent valid tokens for the $j$th decoding step.
Figure 3: Comparing ASAs for continuous and discrete action spaces across 5 environments. For continuous actions, the RVQ tokenization performs best. For discrete actions, SemLang performs best. Each bar gives the average over all tasks in the environment with the full breakdown in Appendix D.
@@ 4.1 Experimental Settings
In all environments, we report the success rate as the fraction of episodes in which the agent completed the language instruction. We use the success criteria provided by each environment. We train a policy per action adapter for each environment and report the generalization performance in the main text. When reporting a single success rate per environment, it is the success averaged between all evaluation episodes containing all tasks. We give the full per-task breakdown for results in Appendix D. CALVIN, Meta-World, HabPick, and BabyAI provide expert demonstrations succeeding at the task. CALVIN has $17.9k$ from humans, Meta-World $22.5k$ from a scripted policy, HabPick $6.7k$ generated from an RL policy, and BabyAI $5k$ from a scripted policy. Full details on the train and evaluation setups per environment are in Appendix B.
@@ 5 Limitations and Conclusion
A limitation of our work is all our analysis is under a single MLLM (LLaVA). Another limitation is that RVQ, the best performing ASA in continuous action spaces, requires collecting demonstrations to train the VQ model. Also, SemLang requires manually describing the actions with language. Finally, our analyses are under a single LoRA training setting. Future analyses can explore different base MLLMs under different training regimes like full LLM finetuning.
@@ B.4 BabyAI
Evaluation: We report performance on the unseen synonyms generalization test, described in Section 4.2 of Carta et al. [40]. We evaluate on 20 episodes per template type, giving 100 total evaluation episodes.
@@ C.3 Training and Architecture Details
Appendix D Per-Task Breakdown
In this section, we show results for each environment by task type. Appendix D shows performance on Language Rearrangement for each of the evaluation datasets. Table 5 shows performance on CALVIN for each of the CALVIN tasks. Table 6 shows performance on BabyAI for each of the BabyAI instruction types. Table 7 shows performance on Meta-World for each of the 45 Meta-World task types.
@@ Appendix D Per-Task Breakdown
open 26 6 21
putnext 7 7 7
pick up seq go to 20 10 25
@@
assembly 95 70 5 60
basketball 85 70 55 100
button-press-topdown 100 90 36 100
button-press-topdown-wall 100 100 57 90
button-press 100 100 85 100
@@
coffee-button 100 100 97 100
coffee-pull 96 40 26 50
coffee-push 75 20 25 80
dial-turn 96 50 36 100
disassemble 55 30 25 50
@@
handle-press 100 100 94 100
handle-pull-side 36 10 6 10
handle-pull 65 20 45 30
lever-pull 55 40 45 40
peg-insert-side 64 70 4 40
peg-unplug-side 54 30 94 90
pick-out-of-hole 45 90 35 30
pick-place 74 20 44 60
pick-place-wall 75 20 35 30
plate-slide 100 60 44 100
@@
plate-slide-back 100 90 14 100
@@
push-back 45 30 15 20
push 64 20 74 90
push-wall 76 40 56 100
reach 34 20 14 70
reach-wall 75 80 75 70
shelf-place 43 20 13 10
soccer 44 0 64 40
stick-push 100 60 5 100
stick-pull 86 50 26 100
sweep-into 65 40 55 70
sweep 85 30 45 70

### v2

@@ 1 Introduction
A key and open challenge in grounding MLLMs, which limits their capabilities in embodied tasks, is the gap between the native output space, natural language, and the action space of embodied agents. This problem is particularly acute in continuous action spaces, where low-level controllers may require a high degree of precision. Across the literature, a number of architectures and ways of handling action spaces have been proposed, but there has not been a systematic study of these designs. Our contributions generalize prior attempts to adapt MLLMs to generate actions through an empirical study on which principles and strategies are necessary to effectively close the gap between the action spaces of MLLMs and embodied agents. We study various grounding re-parameterization strategies, which we refer to as Action Space Adapters (ASAs), across a range of embodiments, action spaces, and environments. In particular, we explore the following types of ASAs: (1) ASAs that directly generate actions from a new prediction policy using the MLLM hidden representations as input; (2) ASAs that reuse the native token space of the MLLM to encode actions; (3) and ASAs that introduce a new token space to encode the actions of the agent while adapting the MLLMs to predict these new tokens.
@@ 3.2.1 Discrete Action Spaces
Non-Semantic Language (Lang): Actions are mapped to language tokens, but instead of semantically meaningful descriptions of the actions as with SemLang, the actions are mapped to sequences of numbers. For example, “pick apple" is represented with the string “5 3". The policy must then output the tokens corresponding to this text to call this pick action. Note that we can pick any text for this mapping and the choice of integers is arbitrary. However, the selected text is not semantically representative of the action.
@@ 3.2.2 Continuous Action Space Adaptors
where $v_{k}\in V$. The codebook $V$ of size $K$ is learned over an offline dataset $\mathcal{D}$ of actions using a VQ-VAE [56] trained with the mean-squared error for action reconstruction and commitment loss. We overwrite $K$ infrequently used tokens from the LLM vocabulary to represent $V$. We defer the full details of this tokenization process to Section C.2.
@@
where $v_{k}^{i}\in V_{i}$ is the $k^{\textrm{th}}$ code from the $i^{\textrm{th}}$ codebook. Such tokenization can be learned using Residual VQ-VAE [RVQ-VAE, 52] on an offline dataset of actions. The actual number of token sequences we can represent is $K^{M}$. Hence, RVQ presents the opportunity to exponentially increase the action space quantization without having to drastically increase the size of the learned individual codebooks.
@@ 3.3 Training
We use LLaVA-1.5-7B [6] as the base MLLM. We finetune the MLLM with interactive (i.e., action-labeled) data to make it more suited for interacting with a embodied and interactive environment.
@@
Reinforcement Learning (RL) from Environment Feedback We can also optionally finetune the MLLM to optimize an environment reward using RL. However, predicting actions in the MLLM token space dramatically increases the number of possible action predictions, with many possible predictions corresponding to no valid action. For example, there are 32,000 tokens in the LLaMA text tokenizer, giving $32,000^{m}$ possible predictions by the model with $m$ tokens per action. This makes exploration difficult in RL as only a small fraction of the possible actions are valid. We therefore use a token filter to restrict the autoregressive sampling to only be from token sequences corresponding to valid actions. The token filter is a function $M(l_{t}^{1},\dots,l_{t}^{j-1})$ that produces a binary mask over all tokens to represent valid tokens for the $j$th decoding step.
Figure 3: Comparing ASAs for continuous and discrete action spaces across 5 environments. For continuous actions, the RVQ tokenization performs best. For discrete actions, SemLang performs best. Each bar gives the average over all tasks in the environment with the full breakdown in Appendix E.
@@ 4.1 Experimental Settings
In all environments, we report the success rate as the fraction of episodes in which the agent completed the language instruction. We use the success criteria provided by each environment. We train a policy per action adapter for each environment and report the generalization performance in the main text. When reporting a single success rate per environment, it is the success averaged between all evaluation episodes containing all tasks. We give the full per-task breakdown for results in Appendix E. CALVIN, Meta-World, HabPick, and BabyAI provide expert demonstrations succeeding at the task. CALVIN has $17.9k$ from humans, Meta-World $22.5k$ from a scripted policy, HabPick $6.7k$ generated from an RL policy, and BabyAI $5k$ from a scripted policy. Full details on the train and evaluation setups per environment are in Appendix B.
@@ 5 Limitations and Conclusion
A limitation of our work is all our analysis is under a single MLLM (LLaVA). Another limitation is that RVQ, the best performing ASA in continuous action spaces, requires collecting demonstrations to train the VQ model. Our analyses are also under only a single LoRA training setting. Future analyses can explore different base MLLMs under different training regimes like full LLM finetuning. While our investigation of ASAs enables connecting a MLLM to various action spaces, the performance of these methods is still subpar for real-robot deployment where high success and safety are critical. MLLMs with the best ASA still struggle on simple environments like BabyAI, only achieving 40% success rate. Further work is needed to improve the performance of these methods for real-world usage. Our investigation also only studies adapting MLLMs through behavioral cloning or on-policy RL. Future work can investigate if the choice of ASA varies when adapting the MLLM with other learning algorithms such as off-policy RL or offline RL.
@@ B.4 BabyAI
Evaluation: We report performance on the unseen synonyms generalization test, described in Section 4.2 of Carta et al. [40]. We evaluate on 200 episodes per template type, giving 1000 total evaluation episodes.
@@ C.3 Training and Architecture Details
Appendix D Qualitative Results
See Figure 8 for qualitative results of results from Figure 3. The RVQ ASA is visualized for Meta-World, CALVIN and Habitat Pick. SemLang is visualized for Language Rearrangement.
(a) The robot successfully picks the stick and pushes the box to the goal position.
(b) The robot only partially lifts the handle and fails to lift it up all the way.
(c) The robot grasps the pink block and lifts it to the goal height.
(d) The robot attempts to grasp the blue block but grasps too high, failing to pick the block.
(e) The robot moves closer to the cleaner bottle with its base and moves the arm to grasp the cleaner. It then returns the end-effector to the resting position to successfully end the task.
(f) The robot correctly finds the lemon in the sink, but the tight sink receptacle results in the arm colliding with the sink and the episode terminating due to excessive collisions.
(g) The robot searches the house, eventually finds the mug and then brings it to the blue table.
(h) The robot picks the strawberry navigates to the counter area, but puts the strawberry on the right counter as opposed to the correct receptacle of the sink.
Figure 8: Qualitative visualizations of successes and failures from the results in Figure 1 of the main paper. The RVQ action space adapter is visualized for Meta-World, CALVIN and Habitat Pick. SemLang is visualized for Language Rearrangement.
Appendix E Per-Task Breakdown
In this section, we show results for each environment by task type. Appendix E shows performance on Language Rearrangement for each of the evaluation datasets. Table 5 shows performance on CALVIN for each of the CALVIN tasks. Table 6 shows performance on BabyAI for each of the BabyAI instruction types. Table 7 shows performance on Meta-World for each of the 45 Meta-World task types.
@@ Appendix E Per-Task Breakdown
open 26 7 21
putnext 8 5 7
pick up seq go to 21 12 22
@@
assembly 100 70 10 60
basketball 90 70 60 100
button-press-topdown 100 90 40 100
button-press-topdown-wall 100 100 60 90
button-press 100 100 70 100
@@
coffee-button 100 100 100 100
coffee-pull 100 40 30 50
coffee-push 80 20 30 80
dial-turn 100 50 40 100
disassemble 60 30 30 50
@@
handle-press 100 100 90 100
handle-pull-side 40 10 10 10
handle-pull 70 20 50 30
lever-pull 60 40 50 40
peg-insert-side 60 70 0 40
peg-unplug-side 50 30 100 90
pick-out-of-hole 50 90 40 30
pick-place 80 20 40 60
pick-place-wall 80 20 40 30
plate-slide 100 60 40 100
@@
plate-slide-back 100 90 10 100
@@
push-back 50 30 20 20
push 60 20 70 90
push-wall 80 40 60 100
reach 30 20 10 70
reach-wall 80 80 80 70
shelf-place 50 20 10 10
soccer 40 0 60 40
stick-push 100 60 10 100
stick-pull 90 50 30 100
sweep-into 70 40 60 70
sweep 90 30 50 70